OpenStudy (anonymous):

Help solving? 2sin3x-1+0

OpenStudy (anonymous):

\[(\sin^{-1} (1/2))/3=x\]

OpenStudy (anonymous):

2 sin (3x) - 1 = 0 ? <- did you mean 2 sin (3x) - 1 = 0 2 sin (3x) = 1 sin (3x) = 1/2 3x = 30, 150, 390 ... degrees so how about: 3x = 180n + (-1)^n * 30 x = 60n + (-1)^n * 10